Let Y = |Z|, where Z ∼ N (0, 1) be a discrete random variable with the following PMF: (i) Find E(Y ) and V ar(Y ). (ii) Find V ar(Y ). (iii) Find the CDF and PDF of Y . 


1
Expert's answer
2022-01-11T17:28:21-0500

The folded normal distribution is a probability distribution related to the normal distribution. Given a normally distributed random variable "X"  with mean "\u03bc"  and variance "\u03c3^2," the random variable "Y = |X|" has a folded normal distribution. 

The mean of the folded distribution is

"\mu_Y=\sigma\sqrt{\dfrac{2}{\pi}}e^{-\mu^2/(2\sigma^2)}+\mu\big[1-2\Phi(-\dfrac{\mu}{\sigma})\big]"

Given "Z\sim N(0, 1), Y=|Z|." Then "\mu=0, \sigma=1"


"\mu_Y=1\sqrt{\dfrac{2}{\pi}}e^{-0}+0\big[1-2\Phi(-0)\big]=\sqrt{\dfrac{2}{\pi}}"

(i)

"E(Y)=\mu_Y=\sqrt{\dfrac{2}{\pi}}"


(ii)

The variance is expressed in terms of the mean:


"\sigma_Y^2=\mu^2+\sigma^2-\mu_Y^2"

Given "Z\sim N(0, 1), Y=|Z|." Then "\mu=0, \sigma=1"

"\sigma_Y^2=0^2+1^2-(\sqrt{\dfrac{2}{\pi}})^2=1-\dfrac{2}{\pi}=\dfrac{\pi-2}{\pi}"

(iii)

The probability density function (PDF) is given by


"f_Y(y; \mu, \sigma^2)=\dfrac{1}{\sqrt{2\pi \sigma^2}}e^{-(y^2-\mu^2)/(2\sigma^2)}"

"+\dfrac{1}{\sqrt{2\pi \sigma^2}}e^{-(y^2+\mu^2)/(2\sigma^2)}"

for "y \u2265 0," and everywhere else.

Given "Z\sim N(0, 1), Y=|Z|." Then


"f_Y(y; 0, 1)=\dfrac{1}{\sqrt{2\pi(1)^2}}e^{-(y^2-0^2)/(2(1)^2)}"

"+\dfrac{1}{\sqrt{2\pi(1)^2}}e^{-(y^2+0^2)/(2(1)^2)}"

"f_Y(y; 0, 1)=\sqrt{\dfrac{2}{\pi}}e^{-y^2/2}"

for "y \u2265 0," and everywhere else.

In our case "Y=|Z|"  follows a half-normal distribution. 

The cumulative distribution function (CDF) is given by


"F_Y(y;0,1)=\displaystyle\int_{0}^y\dfrac{1}{1}\sqrt{\dfrac{2}{\pi}}e^{-x^2/(2\cdot1)}dx"

"F_Y(y;0,1)=\dfrac{2}{\sqrt{\pi}}\displaystyle\int_{0}^{y/\sqrt{2}}e^{-z^2}dz=\text{erf}(\dfrac{y}{\sqrt{2}})"

where erf is the error function.